In 4 years i will be 3 times the age i was 16 years ago
Answer by ankor@dixie-net.com(22740)

In 4 years i will be 3 times the age i was 16 years ago
a + 4 = 3(a-16)
a + 4 = 3a - 48
4 + 48 = 3a - a
52 = 2a
a = 52/2
a = 26 is your present age
: